OpenStudy (didee):

how do I calculate 8 to the power of 2/3? Please show calculations, thanks

Parth (parthkohli):

First look at this thing: \(\Rightarrow\large {8^{2 \over 3} = 8^{1\over 3} \times 8^{1 \over3} } \) Do you agree with this statement?

OpenStudy (didee):

ok, i follow

Parth (parthkohli):

Okay. So, do you know what \(8^{1 \over 3}\) is?

OpenStudy (didee):

4?

Parth (parthkohli):

No. \(8^{1 \over 3} = \sqrt[3]{8}\). Do you know what the cube root of 8 is?

OpenStudy (didee):

Sorry, that's the answer to 8 to the power of 2/3. I'm not sure how to calculate the cube root by hand. We not allowed to use calculators.

Parth (parthkohli):

Cube root of 8 is the number you multiply to itself three times and get 8.

Parth (parthkohli):

Cube root of 1000 is 10, because 10 x 10 x 10 = 1000.

OpenStudy (didee):

ok, so 2^3 = 8

Parth (parthkohli):

Yes, so cube root of 8 is 3. Return back to what I said. \(\Large 8^{2 \over 3} = 8^{1 \over 3} \times {8^{1 \over 3} = 2 \times 2}\)

OpenStudy (didee):

mmm, cube root of 8 is 2? Check your answer. Excellent explanation, thanks

Parth (parthkohli):

Thanks, and no problem.

Parth (parthkohli):

What's 2 x 2? That'd be the final answer.

OpenStudy (didee):

:-) Thanks I get it.

Parth (parthkohli):

You're welcome!

OpenStudy (didee):

Does it work the same for negative numbers? Say (-8) to the power of 1/3? Just for interest sake...

Parth (parthkohli):

Yeah, cube root of -8 is -2.

OpenStudy (didee):

Ok, cool. Thanks again, cheers

Parth (parthkohli):

Cheers! Bring in the wine, eh?

OpenStudy (didee):

lol... ok, ciao